OpenStudy (anonymous):

Please help! Perform the indicated operations in the following problem; (1+i)/(1-2i) leave the answer in the rectangular form a + b i. a. (3+3i)/5 b. (-1+3i)/5 c. 1-3i d. -1-3i

OpenStudy (dumbcow):

multiply top/bottom by conjugate of denominator **conjugate of a+bi = a-bi **

OpenStudy (dumbcow):

\[\frac{1+i}{1-2i}*\frac{1+2i}{1+2i} = \frac{1+i+2i+2i^{2}}{1+2i-2i-4i^{2}}\] ...simplify

OpenStudy (anonymous):

(-1+3i)/5?

OpenStudy (dumbcow):

yes correct

OpenStudy (anonymous):

Thanks so much!

OpenStudy (dumbcow):

your welcome